# Estimating the Reynolds number dependence of the chaotic attractor dimension in two-dimensional Kolmogorov flow

### v1

@@ V.2 Low-dimensional visualizations
The high-dissipation events (yellow/orange/red markers in figure 11) are found in two (possibly three) distinct clusters: one surrounded by lower-dissipation events, labeled ‘1’ in figure 11, and another which sits apart in the visualization and is more diffuse, labeled ‘2’. This second, higher-dissipation region may itself be made up of two sub-classes, though the amount of data in this region makes this difficult to discern. The bursting class can be associated with a subset of latent SVD modes $\ell\in L_{d}$. This is done in figure 12, where the most significant modes are identified by computing the mean projection onto each $\ell$ within the class, $\langle\mathcal{P}_{\ell}(\omega)\rangle_{i}$ ($i=1$ or 2) (the average was taken over all snapshots satisfying $D/D_{lam}>D_{c}$ within manually defined boundaries on the t-SNE plane). The subset of modes with the largest difference $\Delta\langle\mathcal{P}_{\ell}\rangle\coloneq\langle\mathcal{P}_{\ell}\rangle_{1}-\langle\mathcal{P}_{\ell}\rangle_{2}$ are highlighted in red in figure 12 – this defines the subset $L_{d}$.
@@
The physical structures spanned by these modes are revealed in figure 12, by visualizing the decode of the sum of the projection onto these modes $\mathcal{D}(\sum_{\ell\in L_{d}}\mathcal{P}_{\ell}(\omega_{i})u_{\ell})$ for representative snapshots $\omega_{i}$ from the two bursting classes which were indicated in the earlier t-SNE plot (figure 11). Class 1 (the calmer high-dissipation cluster) has a stronger contribution from the SVD mode 0, which was associated with the time-averaged mean flow (see figure 9). The signature of this large-scale mode is clearly visible in the decodes of the subset $L_{d}$ in figure 12. In contrast, there is a weaker projection onto this mode for class 2, the higher dissipation class, which explains its location away from the bulk of the visualization in figure 11. Class 2 instead has a stronger projection onto more of the modes $\ell\in L_{d}$, and also excites more large $\ell$ modes which describe smaller-scale vorticity features.
Figure 12: The two different bursting event classes identified by the t-SNE visualization at $Re=40$. Snapshots are extracted belonging to the bursting event class $i=1,2$, and the average projection onto each mode $\langle\mathcal{P}_{\ell}\rangle_{i}$ of class $i$ is computed. (Top) The difference $\Delta\langle\mathcal{P}_{\ell}\rangle\coloneq\langle\mathcal{P}_{\ell}\rangle_{1}-\langle\mathcal{P}_{\ell}\rangle_{2}$ is shown for each mode $\ell$. A subset of modes $\ell\in L_{d}$ (red bars) appears to be mostly responsible for delineating the two bursting event classes. (Bottom) Two representative snapshots $\omega$ indicated by fuchsia (top panels) and cyan (bottom panels) markers from the two classes of bursting events are shown. These snapshots correspond to the fuchsia and cyan circles and crosses in the left panel of figure 11. The decodes $\mathcal{D}(\sum_{\ell\in L_{d}}\mathcal{P}_{\ell}(\omega)u_{\ell})$ are shown directly to the right of each representative snapshot $\omega$.
@@
Histograms of all pre- and post-bursting events, visualized over the earlier t-SNE map of the state space, are reported in figure 13. There is a clear distinction between pre- and post-bursting events. The plot reveals that that the ‘calmer’ bursting class is not visited in isolation, but is the route of all high-dissipation events back to quiescent dynamics. This is confirmed by inspecting the full trajectories of bursting events on the t-SNE visualization (not shown), which always first hop between the two clusters within the extreme bursting class 2, before moving to the calmer bursting class 1.
The pre-bursting trajectories always visit the t-SNE cluster marked by a red cross in figure 11, indicating that this cluster encapsulates some important pre-bursting vorticity features. To identify the associated distinguishing latent modes, the mean projection onto each mode $\ell$ for the snapshots within this cluster $\langle\mathcal{P}_{\ell}(\omega)\rangle_{b}$ and for all the snapshots in the test dataset sampled from the turbulent attractor $\langle\mathcal{P}_{\ell}(\omega)\rangle_{turb}$ are computed. The differences between these average projections are visualized in the lower panel of figure 13. This computation indicates that the pre-bursting cluster has a weaker projection onto the $\ell=0$ mode representative of typical low-dissipation states. However, a much stronger projection onto the $\ell=1$ mode, which features staggered diagonal stripes of vorticity (see figure 9) similar to unstable directions associated with many exact solutions which resemble the mean flow state (see figure 21), is the distinguishing feature of the pre-bursting dynamical process within this cluster.
We have also confirmed that pre-bursting and post-bursting regions on the visualization are both characterized by small $|\hat{\omega}(1,0)|$, while the bursting regions have a large value for $|\hat{\omega}(0,n)|$ (not shown). This large value in the bursting region is to be expected, as the production rate of Kolmogorov flow can be expressed by $I=-\hat{\omega}_{R}(0,n)/n$. The small value of $|\hat{\omega}(1,0)|$ in the pre-bursting region and the large value of$|\hat{\omega}(0,n)|$ in the bursting region confirms the significant redistribution of energy between these modes before a bursting event, an observation which was also made in the variational approach of [41].
@@
Figure 13: (Top) Two-dimensional visualizations of the co-ordinate vector (12) at $Re=40$ using the t-SNE algorithm [52]. The dotted black lines indicate level curves of the Gaussian kernel density estimates of the representation, with increasingly bold lines indicating increasing local density. The dotted orange line identifies the level curve of $D/D_{lam}=0.15$. The grayscale histogram indicates the distribution of subtrajectories (left) which precede and (right) follow a bursting event, which is defined by a local maximum in the dissipation rate satisfying $D>0.175D_{lam}$. Each subtrajectory is defined by the closest continuous sequence of snapshots satisfying $D<0.15D_{lam}$ to the local maximum, within 10 advective time units. The subtrajectory is also required to start from or finish at a sufficiently quiescent state satisfying $D<0.125D_{lam}$. (Middle) The dissipation of a sample subtrajectory centered on the bursting event is plotted against time in advective time units, colored according to time. The opaque markers denote the pre- and post-bursting subtrajectories extracted, while the transparent markers indicate the neglected bursting event. The same subtrajectories are shown by the connected, similarly colored points in the t-SNE visualization, demonstrating these sample routes (top left) towards and (top right) away from the bursting event. (Bottom) The difference $\Delta\langle\mathcal{P}_{\ell}\rangle\coloneq\langle\mathcal{P}_{\ell}\rangle_{b}-\langle\mathcal{P}_{\ell}\rangle_{turb}$ is shown for each mode $\ell$, where $\langle\mathcal{P}_{\ell}\rangle_{b}$ and $\langle\mathcal{P}_{\ell}\rangle_{turb}$ denote the mean projection of snapshots belonging to the pre-bursting cluster (red cross in figure 11) and all snapshots in the test dataset, respectively.
Figure 14: Two-dimensional visualizations of the co-ordinate vector (12) at (top) $Re=100$ and (bottom) $Re=400$ using the t-SNE algorithm [52], as in figure 11.
@@
At $Re=100$ (top panel of figure 14), many high-dissipation snapshots are spaced across the lower periphery of the two-dimensional visualization. There is also a distinct detached high-dissipation cluster, indicating a distinct class of events which could not be separated cleanly in earlier models [34, 13]. Although the clustering of data is less obvious than at the lower $Re=40$ value, there are still distinct classes of vorticity snapshots that can be identified in this visualization. Some representative snapshots sampled from these clusters are shown in figure 15. This figure indicates that the classes can be delineated based on the appearance of the large scale vortices in the flow – their number, relative location and strength. The sample from the detached high-dissipation cluster (black circle) again indicates an event where the importance of ‘mean flow’ modes (e.g. $\ell=0$ and 1) is diminished.
Figure 15: Representative snapshots from the dense clusters in the t-SNE visualization of the co-ordinate vector (12) at $Re=100$. The position of each snapshot in the visualization is marked by the correspondingly colored marker in figure 14.
@@
Finally, the t-SNE visualization of the most turbulent flow considered at $Re=400$ (bottom panel of figure 14) no longer appears to cluster events based on a dominant dissipation rate. The high dissipation snapshots are distributed fairly uniformly in the latent t-SNE visualization. The exception here are a couple of small pockets of very high dissipation snapshots at the left periphery of the visualization. This behavior is consistent with the prominence of the large vortex pair associated with the condensate, a state which is apparent in most snapshots and hence leads to a more homogeneous latent representation across the dataset. Some representative snapshots of the high density regions of the t-SNE visualization are presented in figure 16, and all show a large scale vortex pair. Distinctions between snapshots are likely based on smaller scale features (e.g. vortex filaments) and the relative location/shape of the dominant vortices.
Figure 16: Representative snapshots from the dense clusters in the t-SNE visualization of the co-ordinate vector (12) at $Re=400$. The position of each snapshot in the visualization is marked by the correspondingly colored marker in figure 14.
@@ V.3 Periodic orbit coverage
Large numbers of solutions have been amassed recently for the Kolmogorov flow studied here [14, 13, 38] at $Re\in\{40,100\}$. These solutions are visualized in figure 17 in terms of their production/dissipation rates, overlayed on the turbulent distribution. These low dimensional visualizations, and others like them, would seem to indicate the ‘dynamical relevance’ [59] of the solutions due to their apparent proximity to the turbulence. However, these visualizations are not nuanced enough to identify the range of events contained in the library of UPOs, and may give a misleading impression as to the relevance of a given solution to the true turbulent dynamics. The t-SNE plots constructed above provide a way to assess the degree to which the range of true vortical events in the flow is covered by the library of solutions.
@@
Figure 19: (Top) Two-dimensional visualizations of the co-ordinate vector (12) at $Re=100$ using the t-SNE algorithm [52]. The dotted black lines indicate level curves of the Gaussian kernel density estimates of the representation, such that increasingly bold lines indicate increasing local density. The dotted orange line indicates the level curve of $D/D_{lam}=0.045$. The points indicate snapshots sampled along the periods of the ‘dynamically relevant’ (light blue) and ‘dynamically irrelevant’ (gray) UPOs in the full library of 252 solutions. 50 snapshots are sampled for each UPO.
@@
A similar visualization is conducted for the UPOs computed at $Re=100$ in figure 19. While the vast majority of UPOs have production/dissipation values consistent with that of the turbulence (see figure 17), the t-SNE map indicates that nearly all UPOs are determined by the autoencoder to be distinct from the turbulent dataset. This is quite striking given the overlap in various diagnostic statistics reported here and elsewhere [37, 14, 38]. The apparent lack of similarity can be diagnosed by considering the projections $\mathcal{P}_{\ell}(\omega)$ onto individual latent directions.
Figure 20: The joint distribution of the latent projections $\mathcal{P}_{0}$ and $\mathcal{P}_{2}$ of snapshots sampled from the periods of UPOs deemed ‘dynamically relevant’ due to their dissipation and energy distributions (blue) and snapshots sampled from the turbulent attractor (gray). The densities are colored on a log scale.
To investigate the cause for the separation of these UPOs from the turbulent attractor, the mean projection onto each mode $\ell$ was computed for all the snapshots sampled from the ‘dynamically relevant’ (as determined naively from the dissipation and energy PDFs) UPOs at $Re=100$, $\langle\mathcal{P}_{\ell}(\omega)\rangle_{UPO}$, and for all snapshots in the test dataset sampled from the turbulent attractor at $Re=100$, $\langle\mathcal{P}_{\ell}(\omega)\rangle_{turb}$. Differences were found in the distributions for large numbers of the latent projections, but the most significant distinction between the turbulence and the UPOs was found in the leading mode $\ell=0$. Joint distributions for both the UPO data and turbulence are reported in figure 20 (the other mode is selected as $\ell=2$ here). This figure highlights the substantial difference in the $\ell=0$ distributions – the UPO/turbulence data overlap only slightly in the tails of the distributions. Recall that in §V the $\ell=0$ mode was decoded and shown to be visually similar to the symmetry-reduced, time-average flow state in the turbulence. The dramatic reduction of importance of this mode should also be considered in the context of the continuation study reported in [38], which showed that many $Re=100$ solutions appear to ‘move away’ from the turbulent attractor as $Re$ increases. The results reported here would suggest that perhaps they are not inside the turbulent attractor, but just ‘close by’ at $Re=100$ – consistent with the continuation results.
@@ VI Conclusion
The combination of the IRMAE approach with a full symmetry reduction yielded a rich latent-space structure, with which individual latent SVD modes could be decoded into physically interpretable structures at all $Re$ values. Two-dimensional visualizations of large turbulent datasets revealed distinct clusters corresponding to different ‘classes’ of vorticity snapshots. At $Re=40$, multiple types of high-dissipation events were identified. Visualizations of finely resolved bursting trajectories highlighted a common route to a bursting event involving a large projection onto a particular dominant latent mode, and a common route back to quiescent flow through one of the high dissipation clusters.
@@
Autoncoders seek to exactly reconstruct their inputs, and hence many of the $O(Re^{1/3})$ dimensions found here may not correspond directly to dynamical variables. For instance, it is known in three-dimensional turbulence that scales smaller than $\sim 5\pi\eta$ (where $\eta$ is the Kolmogorov length) are ‘slaved’ to the larger scales within the turbulent attractor [60, 61]. The equivalent result in two-dimensional turbulence is unknown at present, but having a model learn only the ‘dynamical’ components of vorticity, determined via an appropriate low-pass filter, would presumably simplify the reconstruction problem and allow for a more robust estimate of the $Re$-dependence. We hope to report on this in the near future.

### v2

@@ V.2.1 Bursting classes
The high-dissipation or ‘bursting’ events (yellow/orange/red markers in figure 12) are found in two (possibly three) distinct classes: one surrounded by lower-dissipation events, labeled ‘1’ in figure 12, and another which sits apart in the visualization and is more diffuse, labeled ‘2’. This second, higher-dissipation region may itself be made up of two sub-classes, though the amount of data in this region makes this difficult to discern. The two classes of bursting events can be distinguished by a subset of latent SVD modes $\ell\in L_{d}$. This is done in the right panel of figure 12, where the most significant modes are identified by computing the mean projection onto each $\ell$ within the class, $\langle\mathcal{P}_{\ell}(\omega)\rangle_{i}$ ($i=1$ or 2) (the average was taken over all snapshots satisfying $D/D_{lam}>D_{c}$ within manually defined boundaries on the t-SNE plane). The subset of modes with the largest difference $\Delta\langle\mathcal{P}_{\ell}\rangle\coloneq\langle\mathcal{P}_{\ell}\rangle_{1}-\langle\mathcal{P}_{\ell}\rangle_{2}$ are highlighted in red in the right panel of figure 12 – this defines the subset $L_{d}$.
@@
The physical structures spanned by these modes are revealed in the bottom right panel of figure 12, by visualizing the decode of the sum of the projection onto these modes $\mathcal{D}(\sum_{\ell\in L_{d}}\mathcal{P}_{\ell}(\omega_{i})u_{\ell})$ for representative snapshots $\omega_{i}$ from the two bursting classes which are indicated in the t-SNE plot. Class 1 (the calmer high-dissipation cluster) has a stronger contribution from the SVD mode 0, which was associated with the time-averaged mean flow (see figure 10). The signature of this large-scale mode is clearly visible in the decodes of the subset $L_{d}$ in the bottom right panels of figure 12. In contrast, there is a weaker projection onto this mode for class 2, the higher dissipation class, which explains its location away from the bulk of the visualization in figure 12. Class 2 instead has a stronger projection onto more of the modes $\ell\in L_{d}$, and also excites more large $\ell$ modes which describe smaller-scale vorticity features.
V.2.2 Bursting trajectories
@@
Figure 13: (Top) Two-dimensional visualizations of the co-ordinate vector (12) at $Re=40$ using the t-SNE algorithm [57]. The dotted black lines indicate level curves of the Gaussian kernel density estimates of the representation, with increasingly bold lines indicating increasing local density. The dotted orange line identifies the level curve of $D/D_{lam}=0.15$. The grayscale histogram indicates the distribution of subtrajectories as defined in the text (left) which precede and (right) follow a bursting event. (Middle) The dissipation of a sample subtrajectory centered on the bursting event is plotted against time in advective time units, colored according to time. The opaque markers denote the pre- and post-bursting subtrajectories extracted, while the transparent markers indicate the neglected bursting event. The same subtrajectories are shown by the connected, similarly colored points in the t-SNE visualization, demonstrating these sample routes (top left) towards and (top right) away from the bursting event. (Bottom) The difference $\Delta\langle\mathcal{P}_{\ell}\rangle\coloneq\langle\mathcal{P}_{\ell}\rangle_{b}-\langle\mathcal{P}_{\ell}\rangle_{turb}$ is shown for each mode $\ell$, where $\langle\mathcal{P}_{\ell}\rangle_{b}$ and $\langle\mathcal{P}_{\ell}\rangle_{turb}$ denote the mean projection of snapshots belonging to the pre-bursting cluster (red star in figure 12) and all snapshots in the test dataset, respectively.
Histograms of all pre- and post-bursting events, visualized over the earlier t-SNE map of the state space, are reported in the top left and right panels of figure 13, respectively. There is a clear distinction between pre- and post-bursting events. The plot reveals that that the ‘calmer’ bursting class 1 is not visited in isolation, but is the route of all high-dissipation events back to quiescent dynamics. This is confirmed by inspecting the full trajectories of bursting events on the t-SNE visualization (not shown), which always first hop between the two clusters within the extreme bursting class 2, before moving to the calmer bursting class 1.
The pre-bursting trajectories always visit the t-SNE cluster marked by a red star in figure 12, indicating that this cluster encapsulates some important pre-bursting vorticity features. To identify the associated distinguishing latent modes, the mean projection onto each mode $\ell$ for the snapshots within this cluster $\langle\mathcal{P}_{\ell}(\omega)\rangle_{b}$ and for all the snapshots in the test dataset sampled from the turbulent attractor $\langle\mathcal{P}_{\ell}(\omega)\rangle_{turb}$ are computed. The differences between these average projections are visualized in the lower panel of figure 13. This computation indicates that the pre-bursting cluster has a weaker projection onto the $\ell=0$ mode representative of typical low-dissipation states. However, a much stronger projection onto the $\ell=1$ mode, which features staggered diagonal stripes of vorticity (see figure 10) similar to unstable directions associated with many exact solutions which resemble the mean flow state (see figure 18), is the distinguishing feature of the pre-bursting dynamical process within this cluster.
[…]
V.2.3 Higher $Re$ visualizations
@@
At $Re=100$ (top panel of figure 14), many high-dissipation snapshots are spaced across the lower periphery of the two-dimensional visualization. There is also a distinct detached high-dissipation cluster, indicating a distinct class of events which could not be separated cleanly in earlier models [35, 14]. Although the clustering of data is less obvious than at the lower $Re=40$ value, there are still distinct classes of vorticity snapshots that can be identified in this visualization. Some representative snapshots sampled from these clusters are shown in the top panel of figure 14. This figure indicates that the classes can be delineated based on the appearance of the large scale vortices in the flow – their number, relative location and strength. The sample from the detached high-dissipation cluster (bottom left of the visualization) again indicates an event where the importance of ‘mean flow’ modes (e.g. $\ell=0$ and 1) is diminished.
Figure 14: Two-dimensional visualizations of the co-ordinate vector (12) at (top) $Re=100$ and (bottom) $Re=400$ using the t-SNE algorithm [57], as in figure 12 , with visualizations of representative snapshots from the dense clusters.
@@
Finally, the t-SNE visualization of the most turbulent flow considered at $Re=400$ (bottom panel of figure 14) no longer appears to cluster events based on a dominant dissipation rate. The high dissipation snapshots are distributed fairly uniformly in the latent t-SNE visualization. The exception here are a couple of small pockets of very high dissipation snapshots at the left periphery of the visualization. This behavior is consistent with the prominence of the large vortex pair associated with the condensate, a state which is apparent in most snapshots and hence leads to a more homogeneous latent representation across the dataset. Some representative snapshots of the high density regions of the t-SNE visualization are again presented in the bottom panel of figure 14, and all show a large scale vortex pair. Distinctions between snapshots are likely based on smaller scale features (e.g. vortex filaments) and the relative location/shape of the dominant vortices.
@@ V.3 Periodic orbit coverage
Large numbers of solutions have been amassed recently for the Kolmogorov flow studied here [15, 14, 39] at $Re\in\{40,100\}$. These solutions are visualized in figure 15 in terms of their production/dissipation rates, overlayed on the turbulent distribution. These low dimensional visualizations, and others like them, would seem to indicate the ‘dynamical relevance’ [64] of the solutions due to their apparent proximity to the turbulence. However, these visualizations are not nuanced enough to identify the range of events contained in the library of UPOs, and may give a misleading impression as to the relevance of a given solution to the true turbulent dynamics. The t-SNE plots constructed above provide a way to visualize the degree to which the range of true vortical events in the flow is covered by the library of solutions. As nearby snapshots in the t-SNE visualizations reflect similar snapshots in the truncated latent space, the visualizations effectively reveal whether the UPOs are similar to samples from the chaotic attractor and provide an alternative way to assess dynamical relevance.
V.3.1 $Re=40$ UPOs
@@
V.3.2 $Re=100$ UPOs
A similar visualization is conducted for the UPOs computed at $Re=100$ in the top panel of figure 17. While the vast majority of UPOs have production/dissipation values consistent with that of the turbulence (see figure 15), the t-SNE map indicates that nearly all UPOs are determined by the autoencoder to be distinct from the turbulent dataset. This is quite striking given the overlap in various diagnostic statistics reported here and elsewhere [38, 15, 39]. The apparent lack of similarity can be diagnosed by considering the projections $\mathcal{P}_{\ell}(\omega)$ onto individual latent directions.
Figure 17: (Top) Two-dimensional visualizations of the co-ordinate vector (12) at $Re=100$ using the t-SNE algorithm [57]. The dotted black lines indicate level curves of the Gaussian kernel density estimates of the representation, such that increasingly bold lines indicate increasing local density. The dotted orange line indicates the level curve of $D/D_{lam}=0.045$. The points indicate snapshots sampled along the periods of the ‘dynamically relevant’ (light blue) and ‘dynamically irrelevant’ (gray) UPOs in the full library of 252 solutions. 50 snapshots are sampled for each UPO. (Bottom) The joint distribution of the latent projections $\mathcal{P}_{0}$ and $\mathcal{P}_{2}$ of snapshots sampled from the periods of the ‘dynamically relevant’ UPOs (blue) and snapshots sampled from the turbulent attractor (purple). The densities are colored on a log scale.
To investigate the cause for the separation of these UPOs from the turbulent attractor, the mean projection onto each mode $\ell$ was computed for all the snapshots sampled from the ‘dynamically relevant’ (as determined naively from the dissipation and energy PDFs) UPOs at $Re=100$, $\langle\mathcal{P}_{\ell}(\omega)\rangle_{UPO}$, and for all snapshots in the test dataset sampled from the turbulent attractor at $Re=100$, $\langle\mathcal{P}_{\ell}(\omega)\rangle_{turb}$. Differences were found in the distributions for large numbers of the latent projections, but the most significant distinction between the turbulence and the UPOs was found in the leading mode $\ell=0$. Joint distributions for both the UPO data and turbulence are reported in the bottom panel of figure 17 (the other mode is selected as $\ell=2$ here). This figure highlights the substantial difference in the $\ell=0$ distributions – the UPO/turbulence data overlap only slightly in the tails of the distributions. Recall that in §V the $\ell=0$ mode was decoded and shown to be visually similar to the symmetry-reduced, time-average flow state in the turbulence. The dramatic reduction of importance of this mode should also be considered in the context of the continuation study reported in [39], which showed that many $Re=100$ solutions appear to ‘move away’ from the turbulent attractor as $Re$ increases. The results reported here would suggest that perhaps they are not inside the turbulent attractor, but just ‘close by’ at $Re=100$ – consistent with the continuation results.
@@ VI Conclusion
The combination of the IRMAE approach with a full symmetry reduction yielded a rich latent-space structure, with which individual latent SVD modes could be decoded into physically interpretable structures at all $Re$ values.
Two-dimensional visualizations of large turbulent datasets revealed distinct clusters corresponding to different ‘classes’ of vorticity snapshots. At $Re=40$, multiple types of high-dissipation events were identified. Visualizations of finely resolved bursting trajectories highlighted a common route to a bursting event involving a large projection onto a particular dominant latent mode, and a common route back to quiescent flow through one of the high dissipation clusters.
@@
Autoencoders seek to exactly reconstruct their inputs, and hence many of the $O(Re^{1/3})$ dimensions found here may not correspond directly to dynamical variables. For instance, it is known in three-dimensional turbulence that scales smaller than $\sim 5\pi\eta$ (where $\eta$ is the Kolmogorov length) are ‘slaved’ to the larger scales within the turbulent attractor [65, 66]. The equivalent result in two-dimensional turbulence is unknown at present, but having a model learn only the ‘dynamical’ components of vorticity, determined via an appropriate low-pass filter, would presumably simplify the reconstruction problem and allow for a more robust estimate of the $Re$-dependence. We hope to report on this in the near future.